REQUIRED
Show the effect of the following transactions of Cheshire Stores on the Accounting equation.Use “+” to denote an increase, “–” to denote a decrease and “0” to denote no change tothe elements of the equation. Copy the table in your answer book and follow the example given. Assume that the bank balance is favourable at all times.

Example: Paid the telephone account by cheque, R800

No.

Assets

=

Equity

+

Liabilities

e.g.

–R800

–R800

0

2.1.1

2.1.2

2.1.3

2.1.4

2.1.5

TRANSACTIONS

  1. 2.1.1 Purchased stationery by cheque, R300.

  2. 2.1.2 Bought a computer on credit, R8 000.

  3. 2.1.3 The proprietor cashed a business cheque for personal use, R500.

  4. 2.1.4 Issued a cheque to a creditor to settle an account, R1 800.

  5. 2.1.5 Received interest on investment, R400.

2.2 REQUIRED
Use the information provided below to answer the following questions:

  1. 2.2.1 Calculate the amount payable to the tax authority (e.g. SARS) by Ritz Traders for VAT for
    the year. (5)

  2. 2.2.2 Calculate the selling price of a bag of washing powder including VAT, if Ritz Traders purchased the washing powder at R36 a bag excluding VAT. The washing powder is priced at cost plus 25%. (3)

m

(10)

4

2.2.3 Explain what you understand by the term “Output tax" (Output VAT). (2)

INFORMATION

  • Ritz Traders is a registered VAT vendor. VAT is levied at a rate of 14%.

  • The total sales of Ritz Traders according to the cash register slips for the year amounted to

    R201 000 (including VAT). This amount included zero-rated items that were sold for R30 000.

  • The total purchases for the year excluding VAT amounted to R150 000. This amount included

    R24 000 paid for zero-rated items purchased.
Answer #1

Answer-2.1

Libilities No. Assets Equity R300 2.1.1 -R300 2.1.2 +R8,000 +R8,000 -R500 -R500 2.1.3 2.1.4 -R1,800 -R1,800 2.1.5 +R400 +R400

Answer-2.2.1

Total sale including zero-rated item and VAT is = R201,000

Total sale excluding zero-rated item but including VAT is = R201,000 - R30,000 = R171,000

Total sale excluding VAT = R171,000 / 114 × 100 = R150,000

Hence, output VAT = R171,000 - R150,000 = R21,000

Total purchase excluding VAT but including zero-rated item = R150,000

Total purchase excluding zero-rated item = R150,000 - R24,000 = R126,000

It is assumed that on purchase also 14% VAT is paid.

Hence, input VAT = R126,000 × 14% = R17,640

VAT liability to be paid to tax authority = Output VAT - Input VAT = R21,000 - R17,640 = R3,360

Answer-2.2.2

Cost of washing powder excluding VAT = R36

Input VAT = R36 + 14% = R41.04

Margin on cost price = R41.04 × 14% = R10.26

Selling price excluding VAT = R41.04 + R10.26 = R51.30

Selling price including VAT = R51.30 + 14% = R58.48

Answer-2.2.3

Output VAT: It is a tax charged on the sale of goods. It is charged on the selling price of the goods.

Output VAT is the value added tax that you calculate and charge on your own sales of goods and services if you are registered for VAT.